OpenStudy (anonymous):

2/3 of what number is 16?

OpenStudy (mathstudent55):

Write it as an equation and solve the equation. Let the unknown number be x.

OpenStudy (anonymous):

16/2 = 8 8 * 3 = 24 24 * (2/3) = 16 answer = 24

OpenStudy (mathstudent55):

2/3 of what number is 16? 2/3 of x is 16 2/3 * x = 16 2/3 x = 16 Now multiply both sides by 3/2 to find x.

OpenStudy (elonasushchik):

24